20 ml of a solution containing equal moles of n a 2 c o 3 in any nahco3 required 16 ml of 0.16m hcl solution to the phenolphthalein end point . what volume of a 0.10m h2so4 solution would have been required and methyl orange been used as indicator
1
Expert's answer
2019-06-28T02:46:21-0400

Solution.

"\frac{x(Na2CO3)}{x(NaHCO3)} = \frac{N(HCl)*V(phen)*E(Na2CO3)*2*100*1000*m}{1000*m*N(H2SO4)*(V(met)-2*V(phen)*E(NaHCO3)*100}"




"\frac{x(Na2CO3)}{x(NaHCO3)} = 1:1"

"1 = \frac{N(HCl)*V(phen)*E(Na2CO3)*2}{N(H2SO4)*(V(met)-2*V(phen))*E(NaHCO3)}"

V(met) = 64.3 ml

Answer:

 V(met) = 64.3 ml